OpenStudy (anonymous):

The half-life of a radioactive isotope is the time it takes for a quantity of the isotope to be reduced to half its initial mass. Starting with 210 grams of a radioactive isotope, how much will be left after 5 half-lives?

OpenStudy (anonymous):

@FaiqRaees

OpenStudy (faiqraees):

So at each half life the mass gets halved 1st halflife = 210/2 g 2nd halflife = 105/2 g 3rd halflife = 52.5/2 g 4th halflife = 26.25/2 g 5th halflife = 13.125/2 g = 6.5625 g

OpenStudy (anonymous):

Wait .. What? So how would i answer the question ? Is there a formula ...

OpenStudy (faiqraees):

You can even use the formula to calculate it Amount left = Initial *(1-rate)^halflife

OpenStudy (anonymous):

so would it look like 210 * ( 1 - ? )^5

OpenStudy (anonymous):

I am not sure what goes where the rate is

OpenStudy (faiqraees):

Everytime the mass gets halved Which means the mass gets multiplied by 0.5 So the rate here is 0.5

OpenStudy (anonymous):

Okay so 210 * ( 1 - 0.5) ^5

OpenStudy (faiqraees):

yes

OpenStudy (anonymous):

I got 6.5625 Is that right ?

OpenStudy (faiqraees):

correct

OpenStudy (anonymous):

Oh my goodness ! Thank you ! :)